OpenStudy (anonymous):

The company discovered that it costs $30 to produce 2 widgets, $118 to produce 4 widgets, and $766 to produce 10. Find the cost of producing 6 widgets. Use the quadratic function to solve

OpenStudy (calvin):

set a function as y=ax^2+bx+c. put in the three pairs of numbers. then we got three equations,30=4a+2b+c, 118=16a+4b+c, 766=100a+10b+c. then find out the values of a,b and c. a=8, b=-4, c=6. So the function is y=8x^2-4x+6. put in x=6, we get y=270. So the cost of producing 6 widgets is $270